OpenStudy (anonymous):

2x-10+3x __________ < -5 4

OpenStudy (kendricklamar2014):

Step 1. Simplify: \[\frac{ 5 }{ 4 }x + \frac{ -5 }{ 2 } < -5\] Step 2. Add 5/2 to both sides: \[\frac{ 5 }{ 4 }x + \frac{ -5 }{ 2 } + \frac{ 5 }{ 2 } <-5 + \frac{ 5 }{ 2 }\] \[\frac{ 5 }{ 4 }x < \frac{ -5 }{ 2 }\]

OpenStudy (kendricklamar2014):

Step 3. Multiply both sides by 4/5: \[\frac{ 4 }{ 5 } \times \frac{ 5 }{ 4 }x < \frac{ 4 }{ 5 } \times \frac{ -5 }{ 2 }\] \[x < -2\]

OpenStudy (kendricklamar2014):

and thats the answer..

OpenStudy (shadowlegendx):

@KendrickLamar2014 You simplified some fractions and did some adding behind the scenes. Be sure to explain what you did in each step, and for simplifying fractions, how you did it.

OpenStudy (kendricklamar2014):

Im sorry @ShadowLegendX I was in a hurry. Next time I will remember to do so :)